The maximum strength of the magnetic field in a electromagnetic wave is 900 N/C. What is the maximum strength of the wave's electric field?

900 N/C
3.0 x 10^-6 T
9.0 x 10^-6 T
900 T
Can’t be determined
Answer #1

Answer: Given that maximum strength of magnetic field B(max)=900 N/C

The maximum strength of electric field E(max)-?

We know that E=C*B

where c is velocity of light,=3*10^8 m/s

therefore E(max)= 3*10^8*9*10^2=27*10^10 N/C or V/m

For given options ,cannot be determined.